is approaching.he collects 1 leaf on the first day of the month,2 leaves on the second day and 3 leaves on the third day of the month and so on . on his birthday he finds that he has collected 210 leaves in all. on which day of the month is his birthday
1
Expert's answer
2019-06-05T05:37:54-0400

Let n birthday in digital form, then we have an arithmetic sequence with a difference of 1, where the first term is 1, and the last term is n.

Denote the sum of the first n terms by Sn

the sum of the first n terms of an arithmetic sequence is

Sn = [(a1 + an)/2]*n

According to the question statement,

210 =[(1+n)/2]*n

210 =(n+n2)/2

420 = n+n2

n+n2-420 = 0

n2 +21n -20n -420 = 0

(n-20)*(n+21) = 0

n1 =20 n2 = -21

since the date of the month is a positive date, n = 20.

The 20th day of the month is a birthday.

Answer: 20.